# Dynamic Classification: Leveraging Self-Supervised Classification to Enhance Prediction Performance

### v1

@@ 4.2 Principle of Redundant Prediction
The best model identified in the Evaluation Module will classify both the training and prediction sets. The corresponding subsets within their respective intervals are then obtained, as shown in Table 2.
@@ 2 Training set 2 Model 2 Prediction set 2
In this process, slight perturbations in the classification of the prediction set are inevitable. These may manifest as minor overlaps between prediction intervals, caused by fluctuations in prediction data or minor data entry errors. To resolve this, redundant training is introduced. This technique combines the training data from a target classification interval with data from adjacent intervals. Data from neighboring intervals is selected according to two criteria:
  • The data must be sorted by predicted target values, and the data closest to the target classification is selected.
  • The amount of data selected from neighboring intervals cannot exceed $K/4$, where $K$ is the total number of target training set, $4$ is the empirical value obtained from multiple experiments.
The introduction of redundant training enhances the model’s robustness by expanding the training subset’s data range, incorporating relevant data from neighboring intervals. This helps balance the prediction range of the prediction subset, ensuring that the predicted values within each subset fall within the expected target range, as derived from the $OptimalSegmentationList$ in the Dynamic Classification section. Samples falling outside this range are filtered out, based on the required accuracy, leaving only those predictions that meet the accuracy threshold.
4.3 Excluding Predicted Results
The boundaries of each subset are initially determined by the OptimalSegmentationList, which is obtained through supervised learning from the training set. However, these boundaries can be manually fine-tuned according to accuracy requirements to achieve stable performance and higher precision. This ensures that predictions falling within the valid range are retained, while those deemed inaccurate are removed.
The exclusion process operates as follows: once the predictions are generated, they are compared against the predefined segmentation boundaries. Predictions that fall outside their corresponding subset’s range are flagged as unreliable and subsequently discarded. This ensures that only the most accurate predictions are retained, enhancing the overall reliability of the classification results.
Critically, the exclusion mechanism is designed to guarantee zero missed detections, meaning that all valid predictions are preserved. Within this constraint, adjustments to the segmentation boundaries focus on reducing false positives as much as possible. By dynamically refining these parameters, the exclusion module contributes to a more precise and robust prediction system, allowing for high accuracy with minimal unnecessary exclusions.
@@ 4.4 Algorithm Advantages and Effectiveness
In practical applications, prediction-based models often face two main challenges. First, balancing the impact of samples with lower and higher target values is difficult. When the training set is sampled with a normal distribution, the predictions tend to concentrate around the middle of the value range, and using uniform sampling only slightly mitigates this issue. The second challenge arises when trying to identify inaccurate predictions. After predictions are made, it is necessary to categorize and re-assess them against an accuracy threshold, which often leads to either missed detections or over-exclusion of data.
@@
The Dynamic Classification Algorithm (DCA) effectively addresses these issues by partitioning the data into smaller, well-defined subsets. This segmentation reduces the range within which each model operates, allowing it to focus on narrower, more homogeneous data intervals. As a result, the complexity of predictions is minimized, improving the accuracy of the model. Importantly, this approach ensures that each subset operates independently, reducing errors that might arise from overlapping data distributions or perturbations in predictions.
@@
Without dynamic classification, unsupervised models may fail to properly distinguish between features with similar values, leading to inaccurate predictions, particularly when the feature distributions are broad. Moreover, in traditional prediction structures, re-learning and selection of new models after predictions often exacerbate these issues, making them computationally expensive and inefficient.
@@
The Dynamic Classification Algorithm (DCA) effectively reduces the complexity of predictions by limiting each subset’s data range. By segmenting the data into quantifiable intervals, DCA can precisely filter prediction results, ensuring that only predictions that meet the required accuracy are retained. This method is particularly significant in industrial applications, as it eliminates the need for additional models or complex retraining steps, simplifying the prediction process.
@@
Furthermore, DCA ensures that the model achieves the goal of **zero missed detections** while minimizing **false positives** as much as possible. By carefully adjusting the segmentation boundaries, the algorithm allows for stable, high-accuracy predictions with minimal exclusion. This balancing act is crucial in real-world scenarios, where the precision of the classification process is paramount, and the model must maintain both accuracy and robustness under various operational conditions.
@@
In practical terms, this algorithm has demonstrated its effectiveness by achieving high accuracy with minimal exclusion. Through the use of self-supervised learning and optimal segmentation, the DCA reliably filters predictions, achieving zero missed detections while minimizing false positives. This ability to discard substandard predictions directly based on data characteristics ensures that the model remains robust and accurate, even in real-world, dynamic environments.
@@
The Dynamic Classification Algorithm has shown superior performance in various datasets, offering a promising solution for applications that require high precision, such as industrial forecasting. While the algorithm still presents opportunities for improvement, particularly in automatic parameter tuning and classification efficiency, it has already proven to be a valuable tool for improving prediction accuracy in both small and large-scale classification tasks.
@@ 5.1 Non-Open Source Data
In battery production, consistency is crucial. In production processes involving chemical and sorting stages, cells must undergo discharge testing to verify that their capacity meets the required standards. The test data for this study comes from actual factory production, consisting of three different production batches, and is used to predict cell capacity. The necessary information about these batches has been anonymized and is provided in Table 3.
The data for these three batches follows a normal distribution. Feature extraction primarily relies on characteristics from the middle production stages, such as weight, time, density, and voltage features from the formation and capacity sorting stages. Data preprocessing steps include filtering, temperature compensation, and removing outliers using the interquartile range method. The dynamic classification algorithm can be applied to both uniform and normal distributions, although parameter adjustments may vary.
@@
Total Data Approximately 70,000 Approximately 140,000 Approximately 85,000
@@ 5.1.1 Comparison Plan and Indicator Settings
Due to confidentiality, the specific numerical values can only be presented approximately. The total number of features is approximately 100, of which about 20% originate from intrinsic battery cell characteristics (such as weight, coating density, etc.), while the remaining 80% are derived from data generated during the production process, including voltage, temperature, and capacity data. In the data preprocessing stage, outliers were handled using the interquartile range method. Specifically, 17% of the outliers were filtered from the 22LA dataset, 19% from the 27LA dataset, and 14% from the 13LA dataset. The remaining data underwent normalization and standardization.
Since average accuracy alone is not sufficient to effectively differentiate between models, this study selected more detailed indicators to assess model performance. These include the proportion of samples with prediction accuracy between 99% and 100% relative to the total, and the proportion of samples with prediction accuracy between 99.5% and 100%. The higher the percentage of these two indicators, the greater the reliability and precision of the prediction.
The experiment compared three methods: KMeans model + prediction, GMM model + prediction, and dynamic classification model. All of these models use linear regression as the underlying prediction method, while the direct prediction results (DP) of the linear model are also recorded as comparison data. The training and prediction sets were divided in a 3:7 ratio, using the same features and outlier detection methods for both sets. To ensure reliability, each experiment was verified with two random seeds, generating two sets of random test results.
For subsequent tests, all prediction models will use the same features and parameters, with linear regression serving as the comparison method to further assess performance differences. These detailed indicators allow for a more accurate evaluation of each model’s performance.
@@ 5.1.2 Comparison Analysis
After completing these steps, in order to achieve the goal of zero missed detections and minimal false positives, we must exclude predictions that are inaccurate. To evaluate this step, we introduce comparisons with advanced models like XGBoost, Gaussian classification models, and Random Forest. These models, as advanced methods, are more capable of capturing detailed information, which typically results in better performance.
Since introducing a new model requires a certain amount of data for training, let’s assume we know that some data is predicted inaccurately, and this data will serve as the training set. This assumption is made to better illustrate the advantages of the dynamic classification algorithm in its mechanism. Assume that the data with prediction accuracy below 99% is $J$, and we divide $J$ into 10 parts, supplementing them with randomly selected data that meets the accuracy threshold, ensuring the training set size remains constant at $J$, with consistent features. Each model will use 9 different training sets, where the ratio of inaccurate to accurate data starts at 1:9 and gradually increases until it reaches 9:1. Finally, we will record the missed detection rate, false positive rate, missed detections count, and false positives count. The results from the three models are shown in Fig.4.
Among the experimental results from these three models, all models achieved a good balance with a 4:6 training-to-test data ratio. The corresponding metrics for false positives and false negatives are depicted in the graph. The bar sections represent the miss rate (blue) and overkill rate (red), while the line sections indicate the absolute number of false positives and false negatives. We can observe that as the proportion of non-compliant data in the training set increases, the miss rate significantly decreases, but the overkill rate gradually increases. This result is the combined effect of changes in both the proportion of the training set and the non-compliant data in the prediction set.
In terms of achieving the goal of ’zero missed detections and minimal false positives,’ XGBoost performed the best. When the proportion of non-compliant to compliant data was set at 8:2, XGBoost successfully excluded all non-compliant data, as seen in the low miss rate. However, the false positive rate remained significant, with the number of misclassified samples also notable. These results illustrate the trade-off in balancing the two metrics: even though misclassified instances were minimized, almost half of the remaining data was misidentified. This discrepancy highlights that, in real-world scenarios, it is hard for model to fully achieve the ideal of zero missed detections and minimal false positives, as production data often contain noise and errors that are difficult to discern. The graph visually represents the model’s performance under different training-prediction ratios, showing the progression from a high miss rate to an acceptable level as the proportion of non-compliant data decreases.
@@
(c) Excluding results of Rnadom Forest
Figure 4: Excluding Results of XGBoost with Different Training and Prediction R atios
@@
Under the Dynamic Classification Algorithm (DCA) framework, the model can leverage information derived from effective self-supervised learning to filter out predictions with significant errors. After independent prediction, each predicted result should fall within its designated range. If a sample exceeds this range by more than ±5%, it can be considered inaccurate. If there are specific accuracy requirements, the range can be fine-tuned by adjusting the segmentation values, thus achieving zero missed detections and minimal false positives. This method ensures that the boundaries of each subset are confined within the accuracy limits, ensuring that each prediction meets the desired precision standard.
@@
The range for each prediction is derived from the OptimalSegmentationList, which is initially defined by the initial segmentation values. While the division module automatically generates initial segmentation values based on the distribution of training set samples, manual adjustments of these initial values can provide results that better meet specific accuracy requirements.
@@
In the **13LA dataset**, to directly filter the data with an error of less than 1%, the predictions are excluded using the exclusion module of the dynamic classification to obtain the final prediction. First, the initial division ratio of dynamic classification is manually set as: [0.02, 0.03, 0.1, 0.1, 0.1, 0.1, 0.1, 0.1, 0.1, 0.1, 0.1, 0.1, 0.03, 0.02] Finally, the **LGBT classification model** achieves the best result, and the optimized division ratio, combined with the numerical distribution of the targets in the training set, produces the $OptimalSegmentationList$:[108873.9, 109546.5, 110193.2, 110485.0, 110707.1, 110873.8, 111040.5, 111193.3, 111359.85, 111540.5, 111860.0, 112123.8] The classification error is only **3.39%**. The prediction results are collected and matched back to the division list for exclusion. The exclusion is set to discard the first and last categorization intervals entirely, while the 11 intermediate intervals are expanded by **1.0025 times** each. The non-excluded and excluded results are shown in Table 5, and the number of samples per interval is shown in Table 6. The obtained image results are presented in Fig.5.
Since the error of the dynamic classification is very small, this method allows us to discard **5%** of the data and still achieve **99%** or higher accuracy for all remaining data.
Non-excluded Result 99.9142% 99.9879% 99.7495%
(a) 13LA Non-excluded
(b) 13LA Excluded
(c) 13LA Comparison
Figure 5: Average precision results
@@
From the above results, it is evident that the dynamic classification algorithm (DCA) has several advantages:
  • High Prediction Accuracy: DCA significantly improves prediction accuracy by dividing the data into smaller subsets and making predictions within smaller, controlled ranges, reducing prediction errors.
@@
It is important to note that for DCA to achieve such excellent results, classification errors must remain below around 5%. If classification errors exceed this threshold, unexpected samples may appear in the classification results, distorting the range of classification intervals. This disruption in the classification range would lead to a decline in overall prediction accuracy and make it difficult to achieve the goal of zero missed detections and minimal false positives. Therefore, maintaining classification error below 5% is crucial to the algorithm’s performance.
5.2 Open Source Data
Open-source datasets are sourced entirely from Kaggle. Four non-sequential datasets, representing various fields, were selected to validate the performance of the proposed model. Comparisons were made with the Random Forest model, SVM (with a Gaussian kernel), Linear Regression model, LGBM model, BP Neural Network model, XGBoost model, and the DC algorithm.Four datasets from different domains were chosen for the regression prediction task:
  1. 1.
  2. 2.
    Appliances Energy Prediction: This dataset contains data related to energy usage and environmental conditions within a residential building. It includes measurements of temperature, humidity, and energy consumption of appliances and lights in various rooms, as well as outdoor weather data from a nearby weather station..
  3. 3.
  4. 4.
    Student Performance Dataset: This dataset contains information about student performance in Math and Portuguese language courses, including demographic, family, and school-related features.
Due to the relatively clean data, these datasets were processed within a uniform framework, employing outlier detection through a quartile-based approach and normalization. The training and prediction sets were divided in a 1:1 ratio. The details of the data, including the effects of classification errors and culling features, are presented in Table 7, while Fig.6 visualizes the evaluation results.
The DC algorithm performs admirably across all four datasets, achieving top-tier results, particularly when the classification error is below 5%. However, as the DC algorithm relies on a classification method for subclass prediction, its performance is impacted when the feature correlation is weak, leading to higher classification errors. In cases like this, even after some data is excluded, the results only approach the performance of Random Forest and XGBoost. When analyzing the Estimation Obesity dataset, which benefits from stronger feature support, the classification error is very low. This indicates that the samples in each subset are nearly always accurate, allowing the Optimal Segmentation List to effectively support the culling function. This enables data reorganization without requiring additional models, significantly improving prediction results. Conversely, datasets like Appliances Energy Prediction and Student Performance Prediction, which have higher classification errors, lack strong feature support, making it difficult to surpass the performance of Random Forest and XGBoost. Nonetheless, the DC algorithm shows great potential, particularly when classification errors are low, allowing it to filter out low-accuracy data through hyperparameter design and the culling function, enhancing the final prediction results.

### v2

@@ 4.2 Principle of Redundant Prediction
The best model identified in the Evaluation Module will classify both the training and test sets. The corresponding subsets within their respective intervals are then obtained, as shown in Table 2.
@@ 2 Training set 2 Model 2 Prediction set 2
During this process, it is inevitable that there will be slight disturbances to the classification of the test set, meaning that the distribution of the prediction subset and the training subset in the same subset will be slightly different. Generally speaking, the data distribution of the predicted subset is slightly larger than that of the predicted subset, which is caused by fluctuations in the predicted data or minor data input errors. To solve this problem, redundant training was introduced. This technology combines the training data from the target classification interval with the data from adjacent intervals, enabling it to have a more comprehensive training range. Select the data of adjacent intervals based on two criteria:
  • The data must be sorted by the predicted target values, and the data closest to the target classification is selected.
  • The amount of data selected from neighboring intervals cannot exceed $K/4$, where $K$ is the total number of target training sets and $4$ is the empirical value obtained from multiple experiments.
In this stage, the training subsets within each partition interval will train the same prediction model, such as the linear regression model or the lgmb model, so that each partition interval has the same type but independent model for corresponding predictions. The applying of redundant training enhances the robustness of the model by expanding the data range of the training subset and incorporating relevant data from neighboring intervals. This helps to balance the prediction range of the prediction subset, ensuring that the predicted values within each subset fall within the expected target range, as derived from the $OptimalSegmentationList$ in the Dynamic Classification section. Samples falling outside this range are filtered out, based on the required accuracy, leaving only those predictions that meet the accuracy threshold.
4.3 Principle of Excluding Predicted Results
The boundaries of each subset are initially determined by $OptimalSegmentationList$, which is obtained through self-supervised learning from the training set. However, these boundaries can be manually fine-tuned according to the accuracy requirements to achieve a stable performance and higher precision. This ensures that predictions falling within the valid range are retained, whereas those deemed inaccurate are removed.
The exclusion process operates as follows: Once the predictions are generated, they are compared against predefined segmentation boundaries. Predictions that fall outside their corresponding subset ranges are flagged as unreliable and subsequently discarded. This ensures that only the most accurate predictions are retained, thereby enhancing the overall reliability of the classification results.
Critically, the exclusion mechanism is designed to guarantee zero missed detections, meaning that all valid predictions are preserved. Within this constraint, adjustments to segmentation boundaries focus on reducing false positives as much as possible. By dynamically refining these parameters, the exclusion module contributes to a more precise and robust prediction system, allowing for a high accuracy with minimal unnecessary exclusions.
@@ 4.4 Algorithm Advantages and Effectiveness
In practical applications, prediction-based models face two major challenges. First, it is difficult to balance the impact of samples with lower and higher target values. When the training set is sampled with a normal distribution, the predictions tend to concentrate around the middle of the value range, and using uniform sampling only slightly mitigates this issue. A second challenge arises when attemping to identify inaccurate predictions. After predictions are made, it is necessary to categorize and re-assess them against an accuracy threshold, which often leads to either missed detections or over-exclusion of data.
@@
The DCA effectively addresses these issues by partitioning the data into smaller, well-defined subsets. This segmentation reduces the range within which each model operates, allowing it to focus on narrower and more homogeneous data intervals. Consequently, the complexity of the predictions is minimized,thereby improving the accuracy of the model. Importantly, this approach ensures that each subset operates independently, reducing errors that might arise from overlapping data distributions or perturbations in the predictions.
@@
Without dynamic classification, unsupervised models may fail to properly distinguish between features with similar values, leading to inaccurate predictions, particularly when feature distributions are broad. Moreover, in traditional prediction structures, the re-learning and selection of new models after predictions often exacerbate these issues, making them computationally expensive and inefficient.
@@
The DCA effectively reduces the complexity of predictions by limiting the data range of each subset. By segmenting the data into quantifiable intervals, DCA can precisely filter the prediction results, ensuring that only predictions that meet the required accuracy are retained. This method is particularly significant in industrial applications, because it eliminates the need for additional models or complex retraining steps, thereby simplifying the prediction process.
@@
Furthermore, DCA ensures that the model achieves the goal of zero missed detections while minimizing false positives as much as possible. By carefully adjusting the segmentation boundaries, the algorithm allows for stable and highly, accurate predictions with minimal exclusion. This balancing act is crucial in real-world scenarios, where the precision of the classification process is paramount, and the model must maintain both accuracy and robustness under various operational conditions.
@@
In practical terms, this algorithm has demonstrated its effectiveness by achieving a high accuracy with minimal exclusion. Through the use of self-supervised learning and optimal segmentation, DCA reliably filters predictions, achieving zero missed detections while minimizing false positives. This ability to discard substandard predictions directly based on data characteristics ensures that the model remains robust and accurate, even in real-world, dynamic environments.
@@
The DCA has shown superior performance on various datasets, offering a promising solution for applications that require high precision, such as industrial forecasting. Although the algorithm still presents opportunities for improvement, particularly in automatic parameter tuning and classification efficiency, it has already proven to be a valuable tool for improving prediction accuracy in both small and large-scale classification tasks.
@@ 5.1 Non-Open Source Data
Consistency is crucial in battery production. In production processes involving the chemical and sorting stages, cells must undergo discharge testing to verify that their capacity meets the required standards. The test data for this study come from actual factory production, consisting of three different production batches, and is used to predict the cell capacity. The necessary information about these batches has been anonymized and is provided in Table 3.
@@
Approximately Data 70,000 140,000 85,000
The data for these three batches follow a normal distribution. Feature extraction primarily relies on characteristics from the middle production stages, such as the weight, time, density, and voltage features from the formation and capacity sorting stages. Data preprocessing steps include filtering, temperature compensation, and outlier removal using the interquartile range method. The dynamic classification algorithm can be applied to both uniform and normal distributions, although the parameter adjustments may vary.
@@ 5.1.1 Comparison Plan and Indicator Settings
Owing to confidentiality, only the specific numerical values can be approximated. The total number of features is approximately 100, of which approximately 20% originate from intrinsic battery cell characteristics (such as weight and coating density), whereas the remaining 80% are derived from data generated during the production process, including voltage, temperature, and capacity data. In the data preprocessing stage, outliers are handled using the interquartile range method. Specifically, 17% of the outliers are filtered from the 22LA, 19% from the 27LA dataset, and 14% from the 13LA dataset respectively. The remaining are normalized and standardized. Missing values are removed from the dataset, and no duplicates are found. Although other outlier detection methods are considered, the interquartile range (IQR) method is deemed sufficient for this experiment owing to its robustness. The IQR method is particularly effective in handling skewed distributions and is less sensitive to extreme values, making it a suitable choice for ensuring the stability and accuracy of the model in this case. This method effectively isolates the core data distribution, ensuring that only significant deviations are treated as outliers.
Because the average accuracy alone is not sufficient to effectively differentiate between models, this study selects more detailed indicators to assess model performance. These include the proportion of samples with prediction accuracy between 99% and 100% relative to the total, and the proportion of samples with a prediction accuracy between 99.5% and 100%. The higher the percentage of these two indicators, the greater the reliability and precision of the prediction.
The experiment compared three methods: K-means model + prediction, GMM model + prediction, and the dynamic classification model. All of these models use linear regression as the underlying prediction method, while the direct prediction result (DP) of the linear model is also recorded as comparison data. The training and test sets are divided in a 3:7 ratio, using the same features and outlier detection methods for both sets. To ensure reliability, each experiment is verified using two random seeds, generating two sets of random test results.
For subsequent tests, all prediction models use the same features and parameters, with linear regression serving as the comparison method to further assess performance differences. These detailed indicators allow for a more accurate evaluation of each model performance.
@@ 5.1.2 Comparison Analysis
After completing these steps, to achieve the goal of zero missed detections and minimal false positives, we must exclude inaccurate predictions. To evaluate this step, we introduce comparisons with advanced models such as XGBoost, Gaussian classification models, and Random Forest. As advanced methods, these models are more capable of capturing detailed information, which typically results in better performance.
Because introducing a new model requires a certain amount of data for training, let us assume that some data are predicted inaccurately, and this data will serve as the training set. This assumption makes to better illustrate the advantages of the dynamic classification algorithm in its mechanism. Assume that the data with prediction accuracy below 99% are $J$, and we divide $J$ into 10 parts, supplementing them with randomly select data that meet the accuracy threshold, ensuring that the training set size remains constant at $J$, with consistent features. Each model uses 9 different training sets, where the ratio of inaccurate to accurate data starts at 1:9 and gradually increases until it reached 9:1. Finally, the missed detection rate, false positive rate, missed detection count, and false positives count are recorded. The results of the three models are shown in Fig.4.
@@
(c) Excluding results of Random Forest
Figure 4: Excluding Results of Three Models
Among the experimental results from these three models, all models achieve a good balance with a 4:6 training-to-test data ratio. The corresponding metrics for false positives and false negatives are shown in the graph. The bar sections represent the miss rate (blue) and overkill rate (red), whereas the line sections indicate the absolute number of false positives and false negatives. We can observe that as the proportion of non-compliant data in the training set increases, the miss rate significantly decreases, but the overkill rate gradually increases. This result is the combined effect of changes in both the proportion of the training set and non-compliant data in the test set.
In terms of achieving the goal of ’zero missed detections and minimal false positives’, XGBoost performed the best. When the proportion of non-compliant to compliant data was set at 8:2, XGBoost successfully excluded all non-compliant data, as indicated by the low miss rate. However, the false positive rate remained significant, and the number of misclassified samples was also notable. These results illustrate the trade-off in balancing the two metrics: although misclassified instances were minimized, almost half of the remaining data were misidentified. This discrepancy highlights that, in real-world scenarios, it is difficult for the model to fully achieve the ideal of zero missed detections and minimal false positives, as production data often contain noise and errors that are difficult to discern. The graph visually represents the performance of the model under different training-prediction ratios, the progression from a high miss rate to an acceptable level as the proportion of non-compliant data decreases.
@@
Under the DCA framework, the model leverages information derived from effective self-supervised learning to filter out predictions with significant errors. After the independent prediction, each predicted result falls within its designated range. If a sample exceeds this range by more than ±5%, it is considered inaccurate. If there are specific accuracy requirements, the range can be fine-tuned by adjusting the segmentation values, thus achieving zero missed detections and minimal false positives. This method ensures that the boundaries of each subset are confined within accuracy limits, ensuring that each prediction meets the desired precision standard.
@@
The range for each prediction is derived from OptimalSegmentationList, which is initially defined by the initial segmentation values. While the division module automatically generates initial segmentation values based on the distribution of the training set samples, manual adjustments of these initial values can provide results that better meet specific accuracy requirements.
@@
In the 13LA dataset, to directly filter the data with an error of less than 1%, the predictions were excluded using the exclusion module of the dynamic classification to obtain the final prediction. First, the initial division ratio of dynamic classification was manually set as: [0.02, 0.03, 0.1, 0.1, 0.1, 0.1, 0.1, 0.1, 0.1, 0.1, 0.1, 0.1, 0.03, 0.02] Finally, the LGBT classification model achieved the best result, and the optimized division ratio, combined with the numerical distribution of the targets in the training set, produced the OptimalSegmentationList:[108873.9, 109546.5, 110193.2, 110485.0, 110707.1, 110873.8, 111040.5, 111193.3, 111359.85, 111540.5, 111860.0, 112123.8] The classification error was only 3.39%. Prediction results were collected and matched to the division list for exclusion. The exclusion was set to entirely discard the first and last categorization intervals, while the 11 intermediate intervals were expanded by 1.0025 times each. The non-excluded and excluded results are shown in Table 6, and the number of samples per interval is shown in Table 5. The obtained image results are shown in Fig.5
@@
(a) 13LA Non-excluded
(b) 13LA Excluded
(c) 13LA Comparison
Figure 5: Average precision results
Non-excluded Result 99.9142% 99.9879% 99.7495%
Because the error of the dynamic classification is very small, this method allows us to discard 5% of the data and still achieve 99% or higher accuracy for all the remaining data.
From the above results, it is evident that the DCA has several advantages:
  • High Prediction Accuracy: DCA significantly improves prediction accuracy by dividing the data into smaller subsets and making predictions within smaller, controlled ranges, thereby reducing prediction errors.
@@
It is important to note that for DCA to achieve such excellent results, the classification errors must remain below approxiamtely 5%. If classification errors exceed this threshold, unexpected samples may appear in the classification results, distorting the range of classification intervals. This disruption in the classification range would lead to a decline in the overall prediction accuracy making it difficult to achieve the goal of zero missed detections and minimal false positives. Therefore, maintaining the classification error below 5% is crucial for the performance of the algorithm.
5.2 Open Source Data
The open-source datasets are sourced entirely from Kaggle. Four non-sequential datasets, representing various fields, are selected to validate the performance of the proposed model. Comparisons are made with the Random Forest model, SVM (with a Gaussian kernel), Linear Regression model, LGBM model, BP Neural Network model, XGBoost model, and the DC algorithm.Four datasets from different domains were chosen for the regression prediction task:
  1. 1.
  2. 2.
    Appliances Energy Prediction: This dataset contains data related to energy usage and environmental conditions within a residential building. It includes measurements of the temperature, humidity, and energy consumption of appliances and lights in various rooms, as well as outdoor weather data from a nearby weather station.
  3. 3.
  4. 4.
    Student Performance Dataset: This dataset contains information on student performance in Math and Portuguese language courses, including demographic, family, and school-related features.
Due to the relatively clean data, these datasets were processed within a uniform framework, employing outlier detection through a quartile-based approach and normalization. The training and test sets were divided in a 1:1 ratio. The details of the data, including the effects of classification errors and culling features, are presented in Table 7 and Fig.6 shows the evaluation results.